OpenStudy (zepp):

Note: This is NOT a question; it's a tutorial on linear equation. (I'm re-posting this because it disappeared from OpenStudy somehow..)

OpenStudy (zepp):

\[\large 1)~The~ General~form\] \[Ax + By + C = 0\] (Note here: I used capital A,B and C, because they are constants and they DO NOT represent ANY relevant information about the function. A & B must not be 0. Personally I hate this form.) So yeah… You have nothing! No slope, no x-intercept, no y-intercept. But formulas exist to find all of those. To find the slope, use \[\frac{-A}{B}\]To find the x-intercept, use \[\frac{-C}{A}\]To find the y-intercept, use \[\frac{-C}{B}\]Let me give you an example. Let’s say we have the linear function [5x + 9y + 14 = 0\] The slope would be \[\frac{-A}{B}=\frac{-5}{9}\] The x-intercept would be \[\frac{-C}{A}=\frac{-14}{5}\] The y-intecept would be \[\frac{-C}{B}=\frac{-14}{9}\] \[\large 2)~The~ Slope-Intercept~form\]\[\large y=mx+b\]\[m=slope;~~b=y-intercept\] What’s good with this form? You have the y-intercept and the slope, what you need is the x-intercept. To find the x-intercept, use \[\large \frac{-b}{m}\] An example..: \[\large y = 45x + 8\] The x-intercept would be \[\large \frac{-b}{m}=\frac{-8}{45}\] BEWARE OF TRICKY QUESTIONS! If we have something like \[\large y = 7(14x + 7)\]This is NOT the slope-intercept form, remember to DISTRIBUTE the 7 or any other constant to get it \[\large y = 7(14x + 7)\]\[\large y = 98x + 49\] \[\large 3)~The~ intercept~form\](I love this one)\[\large \frac{x}{a}+\frac{y}{b}=1\]\[a = x-intercept~;~~~~b = y-intercept;~~~~~\frac{-b}{a}=slope.\] See? You have everything here. \[\large 4)~The~ point-slope~form\] This is basically the formula to find a slope by using two points but slightly modified. \[\large y - y_{1} = m(x-x_{1})\]\[m= slope; ~~~~x_{1} ~and~y_{1}: A~point~on~the~function\] This form is *only* used to find the equation of a linear equation when the slope and a point is given. Example, let's say we have a point at (-5,6) and a slope of 3, -5 and 6 are respectively x1 and y1. \[\large y - y_{1} = m(x-x_{1})\]\[\large y - 6 = 3(x--5)\]\[\large y - 6 = 3(x+5)\]Distribute...\[\large y - 6 = 3x+15\]\[\large y = 3x+15+6\]\[\large y = 3x+21\] We have the slope-intercept form! ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ Now.. How to find a equation of a linear function? Situation 1: The slope and the y-intercept are given Situation 2: The slope and the x-intercept are given Situation 3. The slope and a random point of the function are given Situation 4. Two random points of the function are given Situation 5. The x-intercept and the y-intercept are given >>>>>>>>>> Situation 1: The slope and the y-intercept are given I give you... a slope! 8! and the y-intercept! 78! Oh come on.. \[\large y=mx+b\]\[m=slope;~~b=y-intercept\]...\[\large y=8x+78!!!\] Alright alright.. >>>>>>>>>> Situation 2: The slope and the x-intercept are given I give you 19 as slope and 3 as the x-intercept We know that \[\large 2)~The~ Slope-Intercept~form\]\[\large y=mx+b\]\[m=slope;~~b=y-intercept\] To find the x-intercept, use \[\large \frac{-b}{m}\]Plug stuffs in... \[\large \frac{-b}{m}=3\]\[\large \frac{-b}{19}=3\]\[\large \frac{-b}{19}*19=3*19\]\[\large -b=57\]\[\large b=-57\] Back to the situation! Slope 19, y-intercept -57 \[\large y = mx+b\]\[\large y = 19x-57\] >>>>>>>>>> Situation 3. The slope and a random point of the function are given I roll a dice.. I got 2 and 92, so our random point is at (2,92) I roll a dice again... got 3.. D: Point-slope form! \[\large y - y_{1} = m(x-x_{1})\]\[m= slope; ~~~~x_{1} ~and~y_{1}: A~point~on~the~function\] 2 is x1, 92 is y1 \[\large y - y_{1} = m(x-x_{1})\]\[\large y - 92 = 3(x-2)\]\[\large y - 92 = 3(x-2)\]Distribute...\[\large y - 92 = 3x-6\]\[\large y = 3x-6+92\]\[\large y = 3x+86\] >>>>>>>>>> Situation 4. Two random points of the function are given Two random points. (19,31) and (22,16) We know that m(the slope) is Rise/Run, or Delta y/Delta x, well, we are going use it to find the slope.\[m=\frac{16-31}{22-19}=-5\] Okay, our slope is -5. We have a point, and a slope. Point slope form! \[\large y - y_{1} = m(x-x_{1})\]\[m= slope; ~~~~x_{1} ~and~y_{1}: A~point~on~the~function\] 19 is x1, 31 is y1 \[\large y - y_{1} = m(x-x_{1})\]\[\large y - 31 = -5(x-19)\]Distribute...\[\large y - 31 = -5x+95\]\[\large y = -5x+95+31\]\[\large y = -5x+126\] >>>>>>>>>> Situation 5. The x-intercept and the y-intercept are given I give you x-intercept 1 and y-intercept 9 We are done. :) \[\frac{x}{1}+\frac{y}{9}=1\] Credits: @lgbasallote @AccessDenied @amistre64 @shadowfiend \(\underline{\qquad \qquad \qquad \qquad \qquad} \)

OpenStudy (lgbasallote):

dont tell me you rewrote ALL that

OpenStudy (zepp):

No xD

OpenStudy (zepp):

I was smart enough to keep a \(\LaTeX\) version of the tutorial ;)

OpenStudy (anonymous):

everyone is doing tutorials nowadays, someone should make a good one on integration and one on differential calculus where the graph of the derivative is considered. or a tutorial on advanced and complicated questions involving circular functions and logarithms

OpenStudy (zepp):

I shall try that out :D

OpenStudy (anonymous):

Thanks zepp, it will really help me with my maths.

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{gold}{\star \star}\color{orange}{Great \space Job \space Dude \space ;)}}\]

OpenStudy (vishweshshrimali5):

⋆⋆Great Job Dude ;)